OpenStudy (anonymous):

HELP WILL FAN AND MEDAL Solve the equation by factoring. 1/4x^2 = 25

OpenStudy (anonymous):

Do you know the steps to solve an equation by factoring? What do you need to do first?

OpenStudy (anonymous):

I don't know how do this. Please help me.

OpenStudy (anonymous):

First you have to have an equation set equal to zero. So, subtract 25 from both sides: (1/4)x^2-25=0 Do you know how to factor the difference of squares?

OpenStudy (anonymous):

Subtract 25 on both sides... right?

OpenStudy (anonymous):

That was the first step I did. I subtracted 25 from both sides. Now you have a difference of squares. It's the same form as if you had: 4x^2-25 Do you know how to factor that?

OpenStudy (anonymous):

No. :( I'm stupid when it comes to math :'(

OpenStudy (anonymous):

The difference of squares: (a^2)*(x^2)-(b^2) = (ax-b)*(ax+b) So, in the case of your equation, a^2=1/4 and b^2=25 So, a=1/2 and b=5 ((1/2)x-5)*((1/2)x+5) = 0

OpenStudy (anonymous):

When you multiply two things and get zero, one of them has to equal zero. So, since you have (1/2)x-5 and (1/2)x+5, one of those has to equal zero. So, the two equations you have to solve are (1/2)x-5 = 0 and (1/2)x+5 = 0 You will wind up with two answers.